OpenStudy (anonymous):

I know this is a lot, so I'll send $15 through Paypal to the first person who can finish this. Thanks so much!

OpenStudy (anonymous):

Happy Birthday! Or maybe Happy Un–Birthday? Either way, here is a fun birthday trick. Take the month you were born in. Multiply it by 4, add 13, multiply by 25, subtract 200, add the day you were born, multiply by 2, subtract 40, multiply by 50, add the last two digits of your birth year, subtract 10500…. viola! Your birth date! Impress your friends! Now, use your birthday for some more math fun. You can locate graph paper here. 1. Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. 2. Using complete sentences, explain how to convert your birthday function into standard form. 3. Create two additional quadratic functions, g(x) and h(x). A. The function g(x) will open the same direction as f(x), have the same vertex, but will be narrower. B. The function h(x) will open in the opposite direction as f(x), have the same vertex, but will be wider than f(x). Write your functions below and explain in complete sentences why those functions will meet the requirements in the question. 4. Graph your functions. Include your graph below. 5. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. 6. Explain the appropriate translations of your graphs, if you were born two months earlier and 5 days later. Use complete sentences.

OpenStudy (anonymous):

And my birthday's 8/19/97

OpenStudy (texaschic101):

1. 8 x 4 = 32 32 + 13 = 45 45 x 25 = 1125 1125 - 200 = 925 925 + 19 = 944 944 x 2 = 1888 1888 - 40 = 1848 1848 x 50 = 92400 92400 + 97 = 92497 92497 - 10500 = 81997 oh..thats kinda cool, you got your birth month and year.